OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±6) and asymptotes at y=+-3/4x?

OpenStudy (mertsj):

The vertices are on the y axis so the equation has this form: \[\frac{y^2}{a^2}-\frac{x^2}{b^2}=1\]

OpenStudy (mertsj):

a=6

OpenStudy (mertsj):

The asymptotes are: \[y=\frac{\pm a}{b}x\]

OpenStudy (anonymous):

@Mertsj thank you!

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

So \[\frac{6}{b}=\frac{3}{4}\]

OpenStudy (anonymous):

@Mertsj okay what is the next step?

OpenStudy (mertsj):

What is b?

OpenStudy (anonymous):

b=8

OpenStudy (mertsj):

What is a^2?

OpenStudy (anonymous):

a^2=36

OpenStudy (mertsj):

What is b^2?

OpenStudy (anonymous):

b^2=64

OpenStudy (mertsj):

Write the equation.

OpenStudy (anonymous):

so its y^2/36-x^2/64=1?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

@Mertsj thanks again!

OpenStudy (mertsj):

yw